OpenStudy (anonymous):

What is the product of 3 sqaure root 10x * 2 square root 15 x? Simplify if possible. just the answer

OpenStudy (istim):

Well, let's look at those as numerals. Also, this website not operate as an answer mill. Asking for just the answer is rather rude.

OpenStudy (anonymous):

I HAVE ALOT OF QUESTIONS ... I ACTUALLY DO SOME OF THEM THANK YOU VERY MUCH/

OpenStudy (istim):

\[\sqrt[3]{10x}\times \sqrt[2]{15x}\]

OpenStudy (anonymous):

IF YOU DIDN'T WANT TO HELP YOU COULD OF JUST MOVED ON .